# Evaluating NeRFs for 3D Plant Geometry Reconstruction in Field Conditions

### v2

@@ Abstract.
We evaluate different Neural Radiance Fields (NeRFs) techniques for reconstructing (3D) plants in varied environments, from indoor settings to outdoor fields. Traditional techniques often struggle to capture the complex details of plants, which is crucial for botanical and agricultural understanding. We evaluate three scenarios with increasing complexity and compare the results with the point cloud obtained using LiDAR as ground truth data. In the most realistic field scenario, the NeRF models achieve a 74.65% F1 score with 30 minutes of training on the GPU, highlighting the efficiency and accuracy of NeRFs in challenging environments. These findings not only demonstrate the potential of NeRF in detailed and realistic 3D plant modeling but also suggest practical approaches for enhancing the speed and efficiency of the 3D reconstruction process.
Neural Radiance Fields (NeRFs); 3D Geometry Reconstruction; Field Conditions
Figure 1. Using NeRFs for 3D reconstruction of plants in field conditions.
1. Introduction
@@
In recent years, the importance of reconstructing 3D geometry has emerged as a critical area within plant sciences. As global challenges in agriculture and botany become increasingly complex [Pereira, 2017], gaining a detailed understanding of plant structures has become essential. This goes beyond mere visual representation; capturing the intricate details of plant morphology provides valuable insights into their growth, responses to environmental stressors, and physiological processes [Kumar et al., 2012, Paturkar et al., 2021a]. NeRF emerges as a pivotal innovation in the domain of 3D geometry reconstruction.
At its core, NeRF (Neural Radiance Fields) represents a groundbreaking technique that utilizes the capabilities of deep learning to synthesize continuous 3D scenes by modeling the volumetric scene function [Mildenhall et al., 2021]. This approach significantly advances beyond traditional methods such as structure from motion (SfM) [Eltner and Sofia, 2020] and Multi View Stereo (MVS) [Chen et al., 2019], which rely on the analysis of discrete 2D pixels, thus marking a new era in the field of 3D model reconstruction.Unlike its predecessors, NeRF enables the rendering of photorealistic scenes from any viewpoint using a set of 2D images, without necessitating explicit 3D geometry or depth maps. This is primarily due to NeRF’s use of implicit representation models, in contrast to the explicit representations like point clouds in Structure from Motion (SfM) and voxel grids in Multi View Stereo (MVS), which inherently limit the quality of the reconstruction. The implicit representation utilized by NeRF is resolution invariant, allowing for significantly more detailed and granular modeling without the constraints of resolution-dependent methods. This innovation not only enhances the reconstruction of complex scenes but also holds the promise of revolutionizing the 3D modeling of plant structures by providing an unprecedented level of detail and granularity. The versatility and rapid adoption of NeRF as a state-of-the-art technique underscore its significance, with applications ranging from virtual reality [Deng et al., 2022] to architectural reconstructions [Tancik et al., 2022]. Particularly in botany and plant research, NeRF’s ability to capture the finest details offers potential for groundbreaking insights into plant structures, establishing it as a key tool in the advancement of modern botanical studies. However, translating this potential into practical applications presents its own set of obstacles.
1.1. Challenges in 3D Plant Modeling
Contemporary 3D modeling techniques for plant structures face significant challenges when attempting to capture the minute details inherent in plants [Kumar et al., 2012]. The complexity of plants, from their delicate leaf venation [Lu et al., 2009] to intricate branching patterns [Evers, 2011], necessitates models that encompass these specific details. A summary of these is given in Table 1
1.1.1. Technical Challenges
@@
Traditional methodologies like photogrammetry are adept at large-scale reconstructions, but often overlook these subtle intricacies [Paturkar et al., 2019, Wu et al., 2020, Wang et al., 2022]. Utilizing tools such as FARO 3D LiDAR scanners have been adopted for 3D reconstructions of plants like maize [Wu et al., 2020] and tomatoes [Wang et al., 2022]. However, to capture every detail, data from multiple angles is essential. This poses a challenge since scanning from each side using devices like FARO 3D LiDAR scanners is time-consuming, limiting the number of angles that can be efficiently captured. Due to the limited poses, this approach does not scale well to capture minute details in large scenes, and consequently, some desired details might be missed in the final model. Andújar et al. have emphasized that, even with advanced sensors, there are gaps in detailed reconstruction and the preservation of essential information for intricate applications [Andújar et al., 2018]. They also point out that while devices such as the MultiSense S7 from Carnegie Robotics combine lasers, depth cameras, and stereo vision to offer commendable results, the high acquisition costs can be prohibitive. The complexity extends beyond static to dynamic realms.
@@
In addition to the aforementioned challenges, the dynamic nature of flexible objects like plants and environmental dynamism introduce an added dimension of complexity. Plants, unlike static entities, undergo growth, exhibit movement in reaction to environmental stimuli, and demonstrate both diurnal and seasonal variations. The environmental dynamism, coupled with plant behavior, further complicates modeling efforts. Paturkar et al.’s comprehensive investigation underscores that this dynamism inherently complicates the attainment of precise 3D models. Factors such as persistent growth, environmental dynamism, and external perturbations, notably in windy scenarios, jeopardize the consistency of data acquisition during imaging processes [Paturkar et al., 2019, Lu, 2023, Paturkar et al., 2021b]. This complexity necessitates innovative solutions in 3D modeling and data processing.
Furthermore, an array of 3D modeling approaches mandate rigorous [Setkowicz, 2014], an exercise that necessitates vast computational capacities and specialized knowledge. This post-processing phase, being resource and time-intensive, often becomes a constraining factor in projects, particularly when modifications or refinements are imperative based on novel data or feedback. As emphasized in Liénard et al. [2016], discrepancies during the post-processing phase in UAV imagery-centric 3D scene reconstruction might culminate in costly and irreversible repercussions. This necessitates a reevaluation of the cost-benefit ratio inherent in these technologies.
1.1.2. Accessibility Challenges
@@
These factors indicate that the challenges in capturing every facet of plant structures remain, even when employing sophisticated sensors. Financial implications further exacerbate these challenges. Although techniques like terrasential LiDAR scanners provide superior accuracy, their exorbitant costs often render them inaccessible to a significant portion of researchers. Tang et al. delineate that the substantial financial commitment associated with such advanced equipment, combined with the specialized expertise requisite for its operation, significantly limits their adoption within both academic and enthusiast domains [Tang et al., 2022]. Recognizing these barriers, our study proposes an innovative solution.
Table 1. Challenges in 3D Modeling Techniques for Plant Structures
Technical Challenges Accessibility Challenges
Detail limitations in Photogrammetry High acquisition costs
Data acquisition from multiple angles Limited equipment accessibility
Dynamic nature of plants Expertise for operation
Environmental dynamism
Rigorous post-processing
1.2. Objectives and Significance of the Study
This study focuses on a detailed evaluation of advanced Neural Radiance Fields (NeRF) methodologies to assess their applicability and effectiveness in high-resolution 3D modeling of plant structures. Traditional 3D modeling techniques often fall short when it comes to accurately capturing the complex and dynamic morphology of plants [Nguyen et al., 2015]. Our research aims to explore how NeRF can address these limitations, offering a more nuanced and detailed approach to modeling plant structures. Bridging this gap, our study integrates cutting-edge NeRF techniques with mobile technology advancements.
A pivotal aspect of our methodology is the use of mobile technology for data acquisition. By utilizing the widespread availability and sophisticated capabilities of modern smartphones, we aim to make high-quality 3D data collection more accessible and cost-effective. This approach, combined with the NeRF framework’s ability to process a wide variety of image datasets, is expected to revolutionize 3D plant modeling practices, enhancing both inclusivity and efficiency. Next, we delve into the specifics of our research.
@@ 1.2. Objectives and Significance of the Study
An essential part of our study involves a comparative analysis of different NeRF implementations to determine the most effective configurations for specific modeling needs. This includes assessing detail fidelity, computational efficiency, and the ability to adapt to changes in environmental conditions. Such comparative analysis is crucial for establishing benchmarks for NeRF’s current capabilities and for identifying opportunities for future technological improvements. We summarize our contributions as follows:
  1. (1)
@@
    An evaluation of state-of-the-art NeRF techniques across seven different metrics, offering insights for further research.
  3. (3)
    The development of an end-to-end evaluation framework designed to reassess future reconstruction technologies, specifically for 3D modeling in plant sciences.
Furthermore, a comprehensive review of related works and NeRF techniques mentioned in the literature is provided in Table 2, setting the stage for our investigation.
Table 2. Techniques in Literature.
Used original paper implementation.
Used implementation from Framework (NeRFStudio/SDFStudio).
Instant-NGP NeRFacto Mip-NeRF NeRFs Additional Techniques
Towards a Robust Framework for NeRF Evaluation[Azzarelli et al., 2023]
$\times$ $\times$
Analyzing the Internals of Neural Radiance Fields[Radl et al., 2023]
$\times$ $\times$
SteerNeRF: Accelerating NeRF Rendering via Smooth Viewpoint Trajectory[Li et al., 2023]
$\times$ $\times$ NSVF, KiloNeRF, PlenOctree, DIVeR
A Critical Analysis of NeRF-Based 3D Reconstruction[Remondino et al., 2023]
$\times$ $\times$ Tensorf, Mono-Neus, Neus-Facto, MonoSDF, VolSDF, NeuS, UniSurf
Few-Shot Photogrametry: NeRF vs. MVS-SfM in Cultural Heritage Documentation[Balloni et al., 2023]
$\times$ $\times$ $\times$ $\times$
2. Materials and Methods
2.1. Field scenarios
In order to evaluate the efficacy of NeRF techniques across varying complexities and conditions, three scenarios are considered:
  1. (1)
@@ 2.1. Field scenarios
    Single Corn Plant Indoor: This serves as the simplest test case. A solitary corn plant is placed in a controlled indoor environment. The lighting, background, and other environmental factors are kept constant. The objective is to assess the basic capabilities of NeRF in reconstructing an individual plant structure [Jignasu et al., 2023].
  2. (2)
@@
    Multiple Corn Plants Indoor: In this case, more than one corn plant is situated in an indoor setting. The increased complexity due to multiple plants poses a greater challenge for the 3D reconstruction. Inter-plant occlusions and varying plant orientations add an additional layer of complexity.
  3. (3)
@@
    Multiple Corn Plants in a Field with Other Plants: This scenario represents a real-world agricultural field, where corn plants are interspersed with other types of plants. The added complexity due to variable lighting, wind, and other dynamic environmental conditions tests the robustness of the NeRF technology.
2.2. Data Collection
In this experiment, our focus is exclusively on corn plants. We examine three distinct scenarios, elaborated subsequently, spanning from controlled indoor settings to dynamic outdoor ones. In the first scenario, we place a single corn plant in an indoor setting. In the second scenario, we place three corn plants side by side in indoor settings. In the third scenario, in the dynamic outdoor conditions of an experimental field at Iowa State University, we selected a row plot of corn plants planted at approximately 0.2 m distance, approximately at the V12 stage. The leaves between two neighboring plants are overlapping. Transitioning from the specific experimental setups, we now detail the data collection methods utilized in these scenarios.
@@ 2.2. Data Collection
Our training dataset for NeRF is sourced from RGB images and LiDAR data captured using a mobile phone, with the RGB images aiding in the 3D reconstruction of the plants and the LiDAR exclusively for pose capture. For all three scenarios, data is captured using an iPhone 13 Pro featuring 4K resolution. The device is held at a constant height while circling the plant to ensure consistent capture angles. The data collection process utilizes the Polycam app, with a time of approximately 2.5 minutes for scenario 3 (multiple plants in the outdoor setting) and around 1 minute for scenario 1 (single plant in the indoor setting). To establish accurate ground truth, we utilized high-definition terrestrial LiDAR scans using the Faro® Focus S350 Scanner. The scanner boasts an angular resolution of 0.011 degrees, equating to a 1.5 mm point spacing over a 10 m scanning range. With the capacity to acquire point clouds of up to 700 million points (MP) at 1 million points per second. Additionally, the scanner includes a built-in RGB camera that captures 360-degree images once the scanning process is complete.
@@
Both in indoor and outdoor settings, we scan the plants from four (for the single plant) to six (for multiple plants) locations around the plant(s), at a height of 1.5 m and a distance of 1.5 m from the plant(s). To reduce the movement of the leaves during scanning, in indoor settings, we ensure that there is no airflow around the plants, and in outdoor settings, we waited for a suitable time when there was negligible wind flow (August 31, 2023, at 8:30 a.m.). Each scan required approximately 2 minutes and 24 seconds, totaling a capture time of around 18 minutes in outdoor settings, including manually moving the scanner around the plot. The six scans were processed in SCENE® software to add RGB color data to the point clouds, followed by the registration of the clouds by minimizing cloud-to-cloud distance and top view distance. Afterward, we cropped out the area of interest from the registered point cloud, removed duplicate points, and reduced noise using statistical outlier removal based on global and local point-to-point distance distributions. This process resulted in the point cloud having an average resolution of about 7 mm.
This experimental setup enables the NeRF algorithm to work on a range of complexities, from controlled environments to dynamic, real-world conditions.
(a) Frame 1
(b) Frame 2
(c) Frame 3
(d) Frame 4
Figure 2. Raw Images Provided to NeRFs - Scenario I
(a) Frame 1
(b) Frame 2
(c) Frame 3
(d) Frame 4
Figure 3. Raw Images Provided to NeRFs - Scenario II
(a) Frame 1
(b) Frame 2
(c) Frame 3
(d) Frame 4
Figure 4. Raw Images Provided to NeRFs - Scenario III
2.3. Camera Pose, NeRF Training, and Pointcloud Exports
@@
Camera pose estimation is a crucial second step, typically achieved through a Structure from Motion (SfM) pipeline such as colmap. This process is essential for obtaining accurate 3D structures from sequences of images by determining correspondences between feature points and by using sequential matching, especially effective since our dataset comprises video frames.
In our workflow, we incorporate cutting-edge NeRF models for their rapid prototyping capabilities, which are critical in plant sciences to enable large scale reconstructions. Specifically, we employ Instant-NGP [Müller et al., 2022], Nerfacto [Tancik et al., 2023], and TensoRF [Chen et al., 2022], trained on an Nvidia A100 GPU with 80GB of memory. These models are at the forefront of efficiency, with vanilla NeRF requiring approximately 50 times longer to achieve comparable results. Notably, Instant-NGP introduces a small neural network complemented by a multiresolution hash table, optimizing the number of operations required for training and rendering [Müller et al., 2022]. TensoRF, on the other hand, conceptualizes the radiance field as a 4D tensor and applies tensor decomposition to achieve better rendering quality and faster reconstruction times compared to the traditional NeRF approach [Chen et al., 2022].
Nerfacto synthesizes various techniques from multiple contributions in the field, such as the Multilayer Perceptron (MLP) adapted from Instant-NGP, and the Proposal Network Sampler from MipNeRF-360 [Barron et al., 2022], as depicted in the attached image. Post-training, the models are converted into pointclouds with a million points each. The effectiveness of these reconstructions is then rigorously assessed as outlined in the next section.
2.4. Evaluation
2.4.1. Pipline

### v3

# Evaluating Neural Radiance Fields (NeRFs) for 3D Plant Geometry Reconstruction in Field Conditions

@@ Abstract
We evaluate different Neural Radiance Fields (NeRFs) techniques for the 3D reconstruction of plants in varied environments, from indoor settings to outdoor fields. Traditional methods usually fail to capture the complex geometric details of plants, which is crucial for phenotyping and breeding studies. We evaluate the reconstruction fidelity of NeRFs in three scenarios with increasing complexity and compare the results with the point cloud obtained using LiDAR as ground truth. In the most realistic field scenario, the NeRF models achieve a 74.6% F1 score after 30 minutes of training on the GPU, highlighting the efficacy of NeRFs for 3D reconstruction in challenging environments. Additionally, we propose an early stopping technique for NeRF training that almost halves the training time while achieving only a reduction of 7.4% in the average F1 score. This optimization process significantly enhances the speed and efficiency of 3D reconstruction using NeRFs. Our findings demonstrate the potential of NeRFs in detailed and realistic 3D plant reconstruction and suggest practical approaches for enhancing the speed and efficiency of NeRFs in the 3D reconstruction process.
Keywords: Neural Radiance Fields — 3D Reconstruction — Field Conditions
1   Introduction
@@
In recent years, reconstructing 3D geometry has emerged as a critical area within plant sciences. As global challenges in food production become increasingly complex [1], gaining a detailed understanding of plant structures has become essential. This goes beyond mere visual representation; capturing the intricate details of plant geometry provides valuable insights into their growth, responses to environmental stressors, and physiological processes [2, 3]. Consequently, there are several efforts for the 3D reconstruction of plants [4, 5, 6].
Figure 1: NeRFs are proposed as an alternative to traditional TLS scans for 3D plant reconstruction, offering cost-effective and efficient modeling from images captured at multiple angles using a smartphone camera, in contrast to the higher expense and extensive processing time required by TLS for multi-angle scan registration.
@@
One of the most common approaches for 3D reconstruction is photogrammetry, which relies on the analysis of discrete 2D pixels using techniques such as structure from motion (SfM) [7] and multi-view stereo (MVS) [8]. Another direct approach is utilizing LiDAR scanners (such as FARO 3D LiDAR scanner) to capture a dense 3D point cloud of the plants. This approach has been successfully used for the 3D reconstruction of Maize [9] and Tomato plants [10]. Contemporary 3D modeling techniques for plant structures face significant challenges when attempting to capture the minute details inherent in plants [2]. The complexity of plants, from their delicate leaf venation [11] to intricate branching patterns [12], necessitates models that encompass these specific details. Scans from multiple angles are essential to capture every detail, which is challenging since multiple LiDAR scans are time-consuming. Due to the limited poses, this approach does not scale well to capture minute details in large scenes; consequently, some desired details might be missed in the final model. Andújar et al. [13] have emphasized that, even with advanced sensors, there are gaps in detailed reconstruction. They also point out that while devices such as the MultiSense S7 from Carnegie Robotics combine lasers, depth cameras, and stereo vision to offer reasonable results, the high acquisition costs can be prohibitive. At the same time, while photogrammetry is adept at large-scale reconstructions, it often cannot capture subtle details of plants [14, 9, 10].
@@
In addition to the challenges mentioned above, the dynamic nature of flexible objects such as plants and their environment introduces an added complexity. Plants, unlike static entities, undergo growth, exhibit movement in reaction to environmental stimuli such as wind, and demonstrate both diurnal and seasonal variations. The environmental dynamism, coupled with plant behavior, further complicates modeling efforts. Paturkar et al. [14] comprehensive investigation underscores that this dynamism inherently complicates the attainment of precise 3D models. Factors such as persistent growth, environmental dynamism, and external perturbations, notably in windy scenarios, jeopardize the consistency of data acquisition during imaging processes [15, 16]. Liénard et al. [17] highlight that errors in post-processing UAV-based 3D reconstructions can lead to severe, irreversible consequences. This complexity necessitates innovative solutions in 3D modeling and data processing.
One of the most recent approaches for 3D reconstruction is Neural Radiance Fields (NeRFs). At its core, NeRFs utilize deep learning to synthesize continuous 3D scenes by modeling the complete volumetric radiance field [18]. NeRFs enable the rendering of photorealistic scenes from any viewpoint from a neural network trained using a set of 2D images without necessitating explicit 3D geometry or depth maps. NeRFs use implicit representations of the volumetric scene, in contrast to explicit representations such as point clouds in SfM and voxel grids in MVS. The implicit representation utilized by NeRF is resolution invariant, allowing for more detailed and granular modeling without the constraints of resolution-dependent methods. The versatility and rapid adoption of NeRF as a state-of-the-art technique in computer vision and graphics underscore its significance, with applications ranging from virtual reality [19] to architectural reconstructions [20]. Particularly in plant science research, NeRF’s ability to capture fine details offers the potential for deep insights into plant structures and has the potential to be a vital tool in plant phenotyping and breeding (see Figure 1).
@@
These factors indicate that the challenges in capturing detailed plant structures remain, even when employing sophisticated sensors. Financial implications further exacerbate these challenges. Traditional 3D modeling techniques often fall short of accurately capturing the complex 3D structures of plants [21]. Although direct techniques such as LiDAR scanners provide better accuracy, their exorbitant costs often render them inaccessible to many researchers. Tang et al. [22] delineate that the financial commitment associated with such advanced equipment, combined with the specialized expertise requisite for its operation, limits their adoption within academic and enthusiast domains.
@@
In this paper, we perform a detailed evaluation of NeRF methodologies to assess their applicability and effectiveness for high-resolution 3D reconstruction of plant structures. An essential part of our study involves a comparative analysis of different NeRF implementations to determine the most effective framework for specific plant modeling needs. This includes assessing the methods’ fidelity, computational efficiency, and ability to adapt to changes in environmental conditions. Such comparative analysis is crucial for establishing benchmarks for NeRF’s current capabilities and identifying future technological improvement opportunities. Building on this foundation, we introduce an early-stopping algorithm to preemptively terminate the training process, significantly reducing computational cost while retaining model precision. We summarize our contributions as follows:
  1. 1.
@@
    An evaluation of state-of-the-art NeRF techniques across different 2D and 3D metrics, offering insights for further research.
  3. 3.
    An early stopping algorithm to efficiently halt the NeRF training when improvements in model fidelity no longer justify computational costs, ensuring optimal resource use.
  4. 4.
    The development of an end-to-end 3D reconstruction framework using NeRFs designed specifically for the 3D reconstruction of plants.
Our research aims to explore the feasibility of NeRFs for the 3D reconstruction of plants offering an in-depth analysis. A pivotal aspect of our methodology is using low-cost mobile cameras for data acquisition. By utilizing the widespread availability and imaging capabilities of modern smartphones, we can make high-quality image data collection more accessible and cost-effective. This approach, combined with the NeRFs’ ability to process various image datasets for 3D reconstruction, can revolutionize plant reconstruction efforts.
The rest of the paper is arranged as follows. In Section 2, we outline the dataset collection, NeRF implementations, evaluation methods, and the LPIPS-based early-stopping algorithm. In Section 3, we analyze results from single and multiple plant scenarios, both indoors and outdoors, using critical performance metrics. In Section 4, we provide a theoretical discussion on the sampling strategies of different NeRF implementations and examine their impact on performance. We finally conclude in Section 5.
2   Materials and Methods
To evaluate 3D plant reconstruction using NeRFs, we propose a comprehensive methodology encompassing data collection, NeRF implementations, evaluation metrics, and an early stopping algorithm. The overall workflow of the different steps of our framework is shown in Figure 2.
  Reconstruction    Evaluation
Figure 2: Workflow for 3D Reconstruction and Evaluation. The different steps of the above workflow is explained in detailed below.
2.1   Evaluation Scenarios and Data Collection
(a) Scenario I
(b) Scenario II
(c) Scenario III
Figure 3: Example images input to NeRFs for reconstruction across three different scenarios. (a) Scenario I: Indoor single object, (b) Scenario II: Indoor multiple objects, (c) Scenario III: Outdoor scene.
We evaluate NeRFs, examining three distinct scenarios with ground truth data, from controlled indoor to dynamic outdoor environments, and a final testing scenario. The four scenarios are:
  1. 1.
@@ 2.1   Evaluation Scenarios and Data Collection
    Single Corn Plant Indoor: This serves as the simplest test case. A solitary corn plant is placed in a controlled indoor environment. The lighting, background, and other environmental factors are kept constant. The objective is to assess the basic capabilities of NeRF in reconstructing an individual plant structure [23] (see Figure 3(a)).
  2. 2.
@@
    Multiple Corn Plants Indoor: In this case, more than one corn plant is situated in an indoor setting. The increased complexity due to multiple plants poses a greater challenge for the 3D reconstruction. Inter-plant occlusions and varying plant orientations add an additional layer of complexity (see Figure 3(b)).
  3. 3.
@@
    Multiple Corn Plants in a Field with Other Plants: This scenario represents a real-world agricultural field, where corn plants are interspersed with other types of plants. The added complexity due to variable lighting, wind, and other dynamic environmental conditions tests the robustness of the NeRF technology (see Figure 3(c)). We selected a row plot of corn plants planted at approximately 0.2 m distance, approximately at the V12 stage. The leaves between two neighboring plants are overlapping.
  4. 4.
    In-field Test Data: For validating the proposed early stopping methodology, a diverse dataset was assembled, featuring scenarios with Soybean, Anthurium Hookeri, a mixture of plants, Cymbidium Floribundum, and Hydrangea Paniculata (see Figure 8).
@@
Our training dataset for NeRF is sourced from RGB images and LiDAR data captured using a mobile phone, with the RGB images aiding in the 3D reconstruction of the plants and the LiDAR exclusively for pose capture. For all three scenarios, data is captured using an iPhone 13 Pro featuring 4K resolution. The device is held at a constant height while circling the plant to ensure consistent capture angles. The data collection process utilizes the Polycam app [24], with approximately 2.5 minutes for scenario 3 (multiple plants in the outdoor setting) and around 1 minute for scenario 1 (single plant in the indoor setting). To establish accurate ground truth, we utilized high-definition terrestrial LiDAR scans using the Faro® Focus S350 Scanner. The scanner has an angular resolution of 0.011 degrees, equating to a 1.5 mm point spacing over a 10 m scanning range, and the capacity to acquire point clouds of up to 700 million points (MP) at 1 million points per second. Additionally, the scanner includes a built-in RGB camera that captures 360-degree images once the scanning process is complete.
@@
Both in indoor and outdoor settings, we scan the plants from four (for the single plant) to six (for multiple plants) locations around the plant(s) at a height of 1.5 m and a distance of 1.5 m from the plant(s). To reduce the movement of the leaves during scanning, in indoor settings, we ensure that there is no airflow around the plants, and in outdoor settings, we waited for a suitable time when there was negligible wind flow (August 31, 2023, at 8:30 a.m.). Each scan required approximately 2.5 minutes, totaling a capture time of around 18 minutes in outdoor settings, including manually moving the scanner around the plot. The six scans were processed in SCENE® software to add RGB color data to the point clouds, followed by the registration of the clouds by minimizing cloud-to-cloud distance and top view distance. Afterward, we cropped out the area of interest from the registered point cloud, removed duplicate points, and reduced noise using statistical outlier removal based on global and local point-to-point distance distributions. This process resulted in the point cloud having an average resolution of about 7 mm. This experimental setup enables the NeRF algorithm to work on a range of complexities, from controlled environments to dynamic, real-world conditions.
@@
Camera pose estimation is a crucial second step, typically achieved through a Structure from Motion (SfM) pipeline such as COLMAP [25]. This process is essential for obtaining accurate 3D structures from sequences of images by determining correspondences between feature points and by using sequential matching, especially effective since our dataset comprises video frames.
2.2   Neural Radiance Fields (NeRFs)
Neural Radiance Fields (NeRFs) model a scene as a continuous function mapping a 3D position $\mathbf{x}=(x,y,z)$ and a 2D viewing direction $\mathbf{d}=(\theta,\phi)$ to a color $\mathbf{c}=(r,g,b)$ and density $\sigma$. The function is parameterized by a neural network $F_{\theta}$, expressed as:
$(\mathbf{c},\sigma)=F_{\theta}(\mathbf{x},\mathbf{d})$ (1)
Rendering an image involves integrating the color and density along camera rays, a process formalized as:
[…]
We specifically chose Instant-NGP, TensoRF, and NeRFacto to evaluate for plant reconstruction since these implementations are more efficient and achieve comparable results as a vanilla NeRF approximately 50 times faster. Each of these implementations introduces several new features over the vanilla NeRF implementations. Instant-NGP introduces a small neural network complemented by a multiresolution hash table, optimizing the number of operations required for training and rendering [26]. TensoRF, on the other hand, conceptualizes the radiance field as a 4D tensor and applies tensor decomposition to achieve better rendering quality and faster reconstruction times compared to the traditional NeRF approach [27]. NeRFacto combines various techniques such as the Multilayer Perceptron (MLP) adapted from Instant-NGP, and the Proposal Network Sampler from MipNeRF-360 [29]. Apart from these three methods, we also tried the vanilla Mip-NeRF [30]. Unfortunately, Mip-NeRF fails to reconstruct more complicated 3D scenes (such as Scenario II) in our testing. Please refer to the Supplementary section where we provide a table for training (over time) of MipNeRF. We briefly describe the three tested NeRF approaches below.
Instant-NGP: Instant-NGP introduces advancements in NeRFs by focusing on three key improvements: enhanced sampling through occupancy grids, a streamlined neural network architecture, and a multi-resolution hash encoding technique. The hallmark of Instant-NGP is its multi-resolution hash encoding. This approach maps input coordinates to trainable feature vectors stored across multiple resolutions. For each input coordinate, the method hashes surrounding voxel vertices, retrieves and interpolates the corresponding feature vectors, and then inputs these interpolated vectors into the neural network. This process enhances the model’s ability to learn complex geometries and ensures a smoother function due to the trainable nature of the feature vectors. The overall design of Instant-NGP drastically accelerates NeRF training and rendering, enabling near real-time processing capabilities. These enhancements collectively empower Instant-NGP to achieve speedups of up to 1000$\times$. The method also employs multiscale occupancy grids to efficiently bypass empty space and areas beyond dense media during sampling, thereby reducing the computational load. These occupancy grids are dynamically updated based on the evolving understanding of the scene’s geometry, facilitating an increase in sampling efficiency. In parallel, Instant-NGP adopts a compact, fully-fused neural network architecture designed for rapid execution. This network is optimized to operate within a single CUDA kernel, consisting of only four layers with 64 neurons each, resulting in a speed boost—achieving a 5-10 times faster performance than traditional NeRF implementations.
TensoRF: TensoRF improves scene representation by modeling the radiance field as a 4D tensor within a 3D voxel grid, where each voxel is enriched with multi-channel features. This model leverages tensor decomposition to efficiently manage the high-dimensional data, utilizing two key techniques: Canonic Polyadic (CP) and Vector-Matrix (VM) decompositions. CP decomposition simplifies the tensor into rank-one components using compact vectors, reducing the model’s memory footprint. VM decomposition, alternatively, breaks the tensor into compact vector and matrix factors, striking a balance between memory efficiency and detail capture. These enable TensoRF to reduce memory requirements while enhancing rendering quality and accelerating reconstruction times. CP decomposition leads to faster scene reconstruction with improved rendering quality and a smaller model size compared to conventional NeRF approaches. VM decomposition takes this further, offering even better rendering quality and quicker reconstruction, all within a compact model size.
NeRFacto: NeRFacto is an aggregate of techniques optimized for rendering static scenes from real images. The model enhances the NeRF framework by incorporating pose refinement and advanced sampling strategies to improve the fidelity of the scene reconstruction. Pose refinement is critical when initial camera poses are imprecise, which is often the case with mobile capture technologies. NeRFacto refines these poses, thus mitigating artifacts and enhancing detail. The model employs a Piecewise sampler for initial scene sampling, allocating samples to optimize the coverage of both near and distant objects. This is further refined using a Proposal sampler, which focuses on areas that contribute most to the scene’s appearance and is informed by a density function derived from a small, fused MLP with hash encoding. Such a design ensures efficient sampling and better reconstruction. Further explanation and contrast with Instant-NGP is given in the discussion section. The implementations for aforementioned algorithms are taken from the open source project NeRFStudio [28].
Table 1: Recent works comparing the performance of different NeRF techniques for 3D reconstruction applications. Used original implementation. Used implementation in NeRFStudio or SDFStudio.
Paper Instant-NGP NeRFacto TensoRF NeRF Additional Methods
Azzarelli et al. [31] $\times$ $\times$ Mip-NeRF
Radl et al. [32] $\times$ $\times$ Mip-NeRF
Li et al. [33] $\times$ $\times$ NSVF, PlenOctree, KiloNeRF, DIVeR
Remondino et al. [34] $\times$ MonoSDF, VolSDF, NeuS, UniSurf
Balloni et al. [35] $\times$ $\times$ $\times$ -
Ours $\times$ -
There have been several recent works that have compared NeRF approaches for 3D reconstruction. Table 1 summarizes some recent work evaluating different NeRF methodologies. Some of these recent research works also employ additional methods to improve reconstruction fidelity. For example SteerNeRF [33] utilizes neural sparse voxel fields (NSVF) [36], KiloNeRF [37], PlenOctree [38], and DIVeR [39], to obtain a smooth rendering from different viewpoints. NSVF introduces a fast, high-quality, viewpoint-free rendering method using a sparse voxel octree for efficient scene representation. KiloNeRF accelerates NeRF’s rendering by three orders of magnitude using thousands of tiny MLPs, maintaining visual quality with efficient training. PlenOctree uses an Octree data structure to store the Plenoptic function. DIVeR improves upon NeRF by using deterministic estimates for volume rendering, allowing for realistic 3D rendering from few images. Similar to our work, Azzarelli et al. [31] propose a framework for evaluating NeRF methods using Instant-NGP, NeRFacto, and Mip-NeRF, focusing on neural rendering isolation and parametric evaluation. Radl et al. [32] analyze trained vanilla NeRFs, Instant-NGP, NeRFActo, and Mip-NeRF, showing accelerated computations by transforming activation features, reducing computations by 50%.
Remondino et al. [34] analyze image-based 3D reconstruction comparing different NeRFs (including Instant-NGP, NeRFacto, TensoRF, MonoSDF [40], VolSDF [41], NeUS [42], UniSurf [43]) with traditional photogrammetry, highlighting their applicability and performance differences for reconstructing heritage scenes and monuments. Balloni et al. [35] does the same but with using only Instant-NGP. Each of these different NeRF implementations have some advancements over vanilla NeRF. MonoSDF demonstrates that incorporating monocular geometry cues improves the quality of neural implicit surface reconstruction. VolSDF improves the volume rendering of signed distance fields (SDF) using a new density representation. NeuS introduces a bias-free volume rendering method for neural surface reconstruction, outperforming existing techniques in handling complex structures and self-occlusions. UniSurf combines implicit surface models and radiance fields, enhancing 3D reconstruction and novel view synthesis without input masks.
2.3   3D Registration